A wool tapestry is 32 years older than a linen tapestry. Twenty years ago, the wool tapestry was twice as old as the linen tapestry. Find the present age of each.
(I tried drawing chard as we needed to and I wrote now and 20 years ago. i wrote x to the linen tapestry and x+32 to wool tapestry. I wrote x+12 to the 20 years ago part of the wool tapestry and x-20 to the 20 years ago part of linen tapestry. I tried different equations with this but it came out wrong.)
I know that the answer must be 84, 52 but I need the equation. In details please.

Answer by ikleyn(52803)

.
Let W be the wool tapestry age.

Let L be that of the linen tapestry.



Then your first equation is  

W = L + 32          (1)    ("A wool tapestry is 32 years older than a linen tapestry")



Twenty years ago the wool was  (W-20) years,  while linen was (L-20) years.

So, your second equation is 

W-20 = 2*(L-20)     (2)    ("Twenty years ago, the wool tapestry was twice as old as the linen tapestry")



Thus you have this system of two equations

W = L + 32          (1)
W-20 = 2*(L-20)     (2)



If you do understand it, then you understand the key moment in the solution.

The rest is just technique.



From equation (1), substitute this expression  W = L + 32  into the second equation.  You will get

(L+32) - 20 = 2*(L-20).


Now simplify it step by step:

L + 32 - 20  = 2L - 40

L + 12 = 2L - 40

12 + 40 = 2L - L

52 = L.


Answer.  The linen tapestry is 52 years old.  The wood tapestry is 32 years older, 52 + 32 = 84 years.

Solved.

--------------
Is everything clear to you now ?

If not, if you still have questions, do not hesitate to post them to me.
You can do it via the "Thank you" form.
In this case please refer to ID number of the problem, which is 1111793.

==============
There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.